• B.5 The Attaran Corporation manufactures two electrical products: portable air conditioners and portable heaters. The assemPLEASE SHOW HOW TO GRAPH ON EXCEL. I DONT KNOW HOW TO GRAPH IT. please also solve as instructed in problem
Answer #1

Let 'A' and 'H' denote the number of AC and Heaters respectively.

Max Z = 25 A + 15 H
s.t.
3A + 2H <= 240 (wiring time)
2A + 1H <= 140 (drilling time)
A, H >= 0

In Excel, repeat the following table, select it (as shown)

=25*15 3A +2H = 240 2A + 1H <= 140 =(240-3*B3)/2 =140-2*B3 =(240-3*B4)/2 =140-2*B4 =(240-3*B5)/2 =140-2*65 =(240-3*B6)/2 =140

Then go to Insert --> Scatter --> Scatter with smoothed line

We will get this:

-3A +2H<= 240 -2A +1H<= 140 - Isoprofit line 20 40 60 80 100

Then mark the feasibility area (manually inserting 'shapes') and move the isoprofit line away from the origin (mentally).

40,60 -3A +2H<= 240 - 2A +1H<= 140 - Isoprofit line - 20 40 60 80 100

The point at which the isoprofit line leaves the feasibility region is (40, 60) and the optimal solution is:

A = 40
H = 60
Max Z = 25*40 + 15*60 = 1900